# An Enhanced Projection Pursuit Tree Classifier with Visual Methods for Assessing Algorithmic Improvements

### v1

@@ 4 Algorithm comparison
To evaluate both the original algorithm and the proposed modifications, we developed an interactive Shiny app (Chang et al. 2025). The tool includes methods for simulating 2D classification problems, and displays the decision boundaries generated by different algorithms: CART, PPtree, and the proposed extensions (implemented in rpart, PPtreeViz, and PPtreeExt, respectively). The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and is implemented in PPtreeExt.
The app contains three tabs, each tab offers control over data generation and algorithm parameters: Basic-Sim, Sim-Outlier, and MixSim. The first tab simulates well-separated classes, the second includes a cluster of outliers, and the third generates overlapping groups using the MixSim package (Melnykov, Chen, and Maitra 2012). Each tab offers control over data generation and algorithm parameters. The decision boundaries for rpart, PPtree, and the modified PPtree are displayed side by side for visual comparison.
Additional controls, such as the rule partition used in PPtree algorithm and the modified PPtree based on subsetting clases, can be incorporated into the application. These features allow users to examine how changes in rule criteria influence the resulting classification boundaries.
Figure 6 illustrates the use of the interactive application. The first modification (Subsetting clases) is applied to generate the projection, and Rule 1 is used to compute the split point. The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart. Additionally, the resulting decision boundary more accurately partitions the groups in a balanced manner.
Figure 6: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree based on subsetting clases applied to simulated well-separated classes. For PPtree, Rule 1 is used to determine the split points. The modified PPtree produces decision boundaries that are more symmetrically placed between the clusters.
@@
Figure 7 presents a scenario in which Modification 2 (Multiple splits) yields a substantial improvement over the other algorithms. The use of multiple splits enables the small group of outliers from class 2 to be correctly classified.
Figure 7: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree—applied to data in which one class consists of two distinct clusters. The boundaries produced by the modified PPtree with multiple splits better capture the underlying structure of the data compared to the other methods.
Figure 8 compares the performance of the algorithm variants on data simulated from Gaussian mixtures. The user can adjust several parameters of the mixture model: $n$ is the sample size, $BarOmega$ is the desired average overlap between components, $MaxOmega$ specifies the maximum allowable overlap, and $K$ is the number of groups.
Figure 8: Decision boundaries produced by the three algorithms— rpart, PPtree, and the modified PPtree—on data simulated from a Gaussian mixture model.
@@
The interactive application provides insight into how modifications to the algorithm affect model fit. It follows the principle of visualizing the model in the data space (Wickham, Cook, and Hofmann 2015). An additional and important use of the tool has been for debugging and evaluating the algorithmic behavior. By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model—particularly when projections are involved.
This sensitivity is likely to be even greater in higher-dimensional settings. However, in scenarios where class separation occurs along combinations of variables, the use of projected data at each node tends to yield decision boundaries that more closely align with the true data distribution.
5 Performance comparison
@@
This section presents a benchmark data study to evaluate the predictive performance of the proposed method. We compare the extended PPtree models with several classification methods—CART, PPtree, Random Forest, PPforest, and SVM—using 94 benchmark datasets. Table LABEL:tbl-source-tab1 in the Appendix contain the url with the source data while Table LABEL:tbl-data-tab2 summarize the main characteristics of each dataset summarizing the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
@@
For each dataset, two-thirds of the observations are randomly selected for training, and the remaining one-third is used for testing and computing the prediction error. This process is repeated 200 times.
@@
Figure 9: Average test error from 200 training/test splits of benchmark data for four models: rpart, PPtree PPtreeExt modifications 1 and 2, where the PPtree extension performs better than rpart and the original PPtree.
Moreover, in 18 datasets, the modified PPtree outperforms CART, as illustrated in Figure 9, and in only 3 cases it achieves the lowest error among all selected methods, including ensemble-based classifiers. Table LABEL:tbl-datatbl presents an overview of these datasets, which is a subset of Table LABEL:tbl-data-tab2 from the Appendix.
@@
6 Explanation of performance
We would expect that the modified PPtree algorithm outperforms
  • rpart when the best separation between groups is in oblique directions
  • and the original PPtree when class variances are different, or when there are multiple classes.
This can be investigated visually using a tour. There are several strategies for approaching this:
@@ 6 Explanation of performance
    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart and PPtree
@@
      record a sequence of projection bases that shows the boundary created by PPtree
@@
      use the same projection sequence to view the boundaries for the three different methods, possibly use a slice tour to focus on the boundary near the center of the data domain.
Figure 10 show projections of the data20-3 illustrating why PPtreeExt performs better. Group 0 is split on either side of group 1, which means that the original PPtree algorithm would not be able to capture both clusters. We can also see that the separation is mostly on V3 but small amounts of other variables help to improve the separation. This means that rpart will not work quite as well because it will focus on splitting V3.
@@
Figure 10: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective. The group 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal. Secondly, the separation between the two groups is in a combination of variables.
Figure 11 shows some plots illustrating why PPtreeExt outperforms the other methods of the data49-2 data. From Figure 9 we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again. This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging that data20-3. A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure. Plot (a) shows PC 1 vs PC2, with points coloured by the two classes. PC 1 explains 99.8% of the variability in the predictors but it explains nothing of the class structure. This is not uncommon for PCA, where often interesting structure can be found in the smaller components. Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences. Each class appears to have three linear pieces pointing in different directions. This supports why PPtreeExt might work well, because in the combination of variables, the classes are nearly separable with multiple linear cuts. It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do projection pursuit. The optimal projection is shown in plot (c). It can be seen the two classes are mostly separable by a linear combination of variables, but there is no gap between the clusters. The linear combination with so many variables is what causes some difficulties for a simple tree algorithm. The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, means that the PPtreeExt can find better separations than PPtree.
@@
Figure 11: Examining why data49-2 is best classified using PPtreeExt. Because there are 100 predictors, for visualisation it is useful to reduce the dimension using principal component analysis, and then examine the separation between groups. PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a). PC2 and PC3 (b) have an interesting difference between the two classes which shows why multiple linear splits might be useful. A guided projection pursuit on the first 26 principal components (c) shows that these classes could be separated with a linear combination. This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
Figure 12 has plots illustrating why the PPtreeExt performs better on data69-1. From Figure 9 we see that all of the PPtree versions perform better than an rpart fit. This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl). This is sufficiently small to examine using a tour, and we initially use a guided tour with the LDA index. This produced the projection shown in plot (a) where we see the primary difference between the three classes. These are each similar size, but oriented in such a way that in a 2D projection they form a triangle. Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively. The boundary points are generated by simulating a random sample of observations between (-3, 3) in each dimension. The slice cuts through to centre of the box. We can see that PPtreeExt essentially fits the data shape neatly but the rpart boundary is very messy. This data needs oblique cuts, which PPtreeExt provides, and the modified algorithm handles the variance differences better than the original.
@@
Figure 12: Examining why data49-2 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt fitted model and (c) boundary produced by rpart fitted model.
@@ 7 Discussion
This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt and evaluated using a purpose-built interactive web application and real data. The first modification involves subsetting classes when calculating the projection direction at each node, while the second allows for multiple splits per class based on individual entropy evaluations between candidate splits. These changes result in a more flexible classification method that leverages linear combinations of variables. The web application also demonstrates how easily diagnostic tools can be constructed to explore and understand the behavior of new algorithms.
The benchmark data study showed that PPtreeExt predictive performance is better than, CART and PPtree in 18 data out of 94. The modified PPtree algorithm outperforms CART in cases when the best separation between groups is in oblique directions and outperforms the original PPtree when class variances are different, or when there are multiple classes.
@@
Interpretability is increasingly important, because predictions from algorithms may need to be defended in practice. We have shown how to examine the data and the fitted model to understand why the PPtree extension performs better than other tree methods on some data sets. This approach is generally useful for assessing and diagnosing model fits.
@@
The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva, Cook, and Lee 2025). Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree. Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries. This can be accomplished using the current PPforest package.
8 Appendix
Table 3: Source data information
Name URL
data12 https://archive.ics.uci.edu/dataset/518/speaker+accent+recognition
data13-1 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data13-2 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data20-1 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-2 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-3 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data21 https://archive.ics.uci.edu/dataset/537/cervical+cancer+behavior+risk
data22-1 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data22-2 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data26 https://archive.ics.uci.edu/dataset/308/gas+sensor+array+under+flow+modulation
data27-a11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a21 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data29 https://archive.ics.uci.edu/dataset/282/lsvt+voice+rehabilitation
data32 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data34 https://archive.ics.uci.edu/dataset/267/banknote+authentication
data36 https://archive.ics.uci.edu/dataset/236/seeds
data37-1 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-2 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-3 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-4 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-5 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-8 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-14 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data38 https://archive.ics.uci.edu/dataset/230/planning+relax
data39-1 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
Name URL
data39-2 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-3 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-4 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-5 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-6 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-7 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-8 https://archive.ics.uci.edu/dataset/225/gas+sensor+array+drift+dataset
data39-9 https://archive.ics.uci.edu/dataset/226/gas+sensor+array+drift+dataset
data41-1 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-2 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-3 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data42 https://archive.ics.uci.edu/dataset/192/breast+tissue
data44 https://archive.ics.uci.edu/dataset/181/libras+movement
data49-2 https://archive.ics.uci.edu/dataset/167/hill+valley
data49-3 https://archive.ics.uci.edu/dataset/168/hill+valley
data49-4 https://archive.ics.uci.edu/dataset/169/hill+valley
data50 https://archive.ics.uci.edu/dataset/161/mammographic+mass
data52-1 https://archive.ics.uci.edu/dataset/96/spectf+heart
data52-2 https://archive.ics.uci.edu/dataset/96/spectf+heart
data53 https://archive.ics.uci.edu/dataset/139/synthetic+control+chart+time+series
data55-1 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data55-2 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data56-1 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data56-2 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data57 https://archive.ics.uci.edu/dataset/39/ecoli
data61-1 https://archive.ics.uci.edu/dataset/54/isolet
data62 https://archive.ics.uci.edu/dataset/74/musk+version+1
data63 https://archive.ics.uci.edu/dataset/75/musk+version+2
data64-1 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data64-2 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data65 https://archive.ics.uci.edu/dataset/15/breast+cancer+wisconsin+original
data66 https://archive.ics.uci.edu/dataset/62/lung+cancer
data67 https://archive.ics.uci.edu/dataset/59/letter+recognition
data68-1 https://archive.ics.uci.edu/dataset/50/image+segmentation
data68-2 https://archive.ics.uci.edu/dataset/50/image+segmentation
data69-1 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data69-2 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data70 https://archive.ics.uci.edu/dataset/42/glass+identification
data71-1 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
Name URL
data71-2 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
data72-1 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-2 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-3 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-4 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-5 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-6 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-7 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data73 https://archive.ics.uci.edu/dataset/151/connectionist+bench+sonar+mines+vs+rocks
data74 https://archive.ics.uci.edu/dataset/152/connectionist+bench+vowel+recognition+deterding+data
NCI60 https://github.com/natydasilva/PPforest/blob/master/data/NCI60.rda
crab https://github.com/natydasilva/PPforest/blob/master/data/crab.rda
fishcatch https://github.com/natydasilva/PPforest/blob/master/data/fishcatch.rda
glass https://github.com/natydasilva/PPforest/blob/master/data/glass.rda
image https://github.com/natydasilva/PPforest/blob/master/data/image.rda
leukemia https://github.com/natydasilva/PPforest/blob/master/data/image.rda
lymphoma https://github.com/natydasilva/PPforest/blob/master/data/lymphoma.rda
olive https://github.com/natydasilva/PPforest/blob/master/data/olive.rda
parkinson https://github.com/natydasilva/PPforest/blob/master/data/parkinson.rda
wine https://github.com/natydasilva/PPforest/blob/master/data/wine.rda
Table 6: Overview of benchmark datasets
dataset Cases Predictors Groups Imbalance Correlation
data12 329 12 6 0.41 0.38
data13-1 122 13 2 0.03 0.41
data13-2 121 13 2 0.29 0.37
data20-1 2665 5 2 0.27 0.81
data20-2 9752 5 2 0.58 0.29
data20-3 8143 5 2 0.58 0.45
data21 72 19 2 0.42 0.22
data22-1 258 5 4 0.25 0.12
data22-2 145 5 4 0.14 0.18
data26 58 435 4 0.21 0.64
data27-a11 1747 18 5 0.38 0.24
data27-a21 1264 18 5 0.35 0.22
data27-a31 1834 18 5 0.28 0.27
data27-a32 1830 32 5 0.28 0.13
data27-b11 1072 18 5 0.31 0.27
data27-b12 1069 32 5 0.32 0.13
data27-b31 1424 18 5 0.21 0.24
data27-b32 1420 32 5 0.21 0.11
data27-c11 1111 18 5 0.14 0.27
data27-c12 1107 32 5 0.13 0.11
data27-c31 1448 18 5 0.17 0.27
data27-c32 1444 32 5 0.17 0.11
data29 126 310 2 0.33 0.28
data32 403 6 4 0.20 0.11
data34 1372 4 2 0.11 0.43
data36 210 7 3 0.00 0.61
data37-1 29101 53 5 0.23 0.17
data37-14 769 53 2 0.52 0.16
data37-2 14132 53 3 0.62 0.18
data37-3 11835 53 4 0.34 0.16
data37-4 16357 53 5 0.29 0.15
data37-5 17560 53 5 0.21 0.16
dataset Cases Predictors Groups Imbalance Correlation
data37-8 22990 53 9 0.27 0.17
data38 182 12 2 0.43 0.21
data39-1 445 128 6 0.15 0.55
data39-2 1239 128 5 0.35 0.40
data39-3 1586 128 5 0.17 0.56
data39-4 161 128 5 0.32 0.43
data39-5 197 128 5 0.22 0.57
data39-6 2300 128 6 0.25 0.61
data39-7 3613 128 6 0.11 0.68
data39-8 294 128 6 0.43 0.57
data39-9 470 128 6 0.10 0.42
data41-1 5456 2 4 0.34 0.03
data41-2 5456 4 4 0.34 0.15
data41-3 5456 24 4 0.34 0.17
data42 106 9 6 0.08 0.51
data44 360 90 15 0.00 0.27
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data49-4 606 100 2 0.01 0.99
data50 830 5 2 0.03 0.23
data52-1 80 44 2 0.00 0.25
data52-2 187 44 2 0.84 0.29
data53 600 60 6 0.00 0.46
data55-1 3823 64 10 0.00 0.12
data55-2 1797 64 10 0.01 0.12
data56-1 7494 16 10 0.01 0.27
data56-2 3498 16 10 0.01 0.28
data57 327 7 5 0.38 0.23
data61-1 6238 617 26 0.00 0.17
data62 476 166 2 0.13 0.26
data63 6598 166 2 0.69 0.27
data64-1 4435 36 6 0.15 0.49
data64-2 2000 36 6 0.13 0.50
data65 683 9 2 0.30 0.60
data66 27 56 3 0.07 0.20
dataset Cases Predictors Groups Imbalance Correlation
data67 20000 16 26 0.00 0.18
data68-1 210 19 7 0.00 0.30
data68-2 2100 19 7 0.00 0.28
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
data70 214 9 6 0.31 0.23
data71-1 43500 9 7 0.78 0.19
data71-2 14494 9 5 0.79 0.20
data72-1 94 18 4 0.09 0.39
data72-2 94 18 4 0.05 0.40
data72-3 94 18 4 0.10 0.42
data72-4 94 18 4 0.12 0.42
data72-5 94 18 4 0.10 0.49
data72-6 94 18 4 0.06 0.40
data72-7 94 18 4 0.06 0.44
data73 208 60 2 0.07 0.23
data74 990 12 11 0.00 0.22
NCI60 56 31 7 0.05 0.56
crab 200 6 4 0.00 0.95
fishcatch 159 7 7 0.31 0.46
glass 214 10 6 0.31 0.23
image 2310 19 7 0.00 0.28
leukemia 72 41 3 0.40 0.44
lymphoma 80 51 3 0.41 0.75
olive 572 9 9 0.32 0.35
parkinson 195 23 2 0.51 0.50
wine 178 14 3 0.13 0.30

### v2

@@ 4.1 Exploring the algorithm changes interactively
The Shiny app includes methods for simulating 2D classification problems, and displays the decision boundaries generated by three different algorithms: CART, PPtree, and the proposed extensions as implemented in the rpart, PPtreeViz, and PPtreeExt packages, respectively.
There are three data generation methods, “Basic-Sim”, “Sim-Outlier”, and “MixSim”. Basic-Sim simulates classes with separations between clusters on linear combinations of variables, by user-specified correlations between variables for each cluster. Sim-Outlier uses the same approach with the addition of a cluster of outliers corresponding to one of the classes. These two methods allow only for three-cluster data. The user can control the number of observations in each class, the class means, and the Sim-Outlier mean and shape of the outlier groups. MixSim uses a Gaussian mixture model using the MixSim package (Melnykov et al. 2012) to simulate class clusters, with any number of classes. While all classes have equal sample size, the user can choose this and also specify values for the MixSim algorithm. Figure 5, Figure 6, Figure 7 show the interface for the three different data-generating methods. Clicking “Ok” will simulate new samples.
The decision boundaries for rpart, PPtree, and the modified PPtree (implemented in PPtreeExt) are displayed side by side for visual comparison. The user can choose which modifications of PPtree to show. For the other algorithms, only the default parameter choices are used.
Figure 5: Comparing the fits for three-class data with a strong correlation between variables. Modification 1 (different subsetting rules) of PPtree is shown. The modification produces lower error and decision boundaries that are more symmetrically placed between the class clusters.
In Figure 5, three-class data with a strong correlation between variables is simulated. Modification 1 (different subsetting rules) is applied, and rule 1 (Table LABEL:tbl-rules) is used to compute the split point. The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart. The change in boundaries is subtle, but more satisfying. Between classes S (violet squares) and T (orange triangles), the boundary better matches the separation - it is not pulled away from this split by using LDA on the super-group (T and C).
Figure 6: Comparing the fits for data where one class has two distinct clusters. Modification 2 (multiple splits) of PPtree is used. The modification produces lower error and better decision boundaries.
@@
Figure 6 shows data where one class has two separated clusters, as simulated by Sim-Outlier. The PPtree extension used is modification 2 (multiple splits). It yields a substantial improvement over the other algorithms, both in lower error and more sensible boundaries. The use of multiple splits enables the small group of outliers from class T (orange triangles) to be better classified.
The data in Figure 7 has 5 clusters of different shapes. Modification 2 (multiple splits) is examined. Again, we see that the modified algorithm outperforms PPtree and rpart, in both lower error and better boundaries.
@@
The interactive application provides insight into how modifications to the algorithm affect model fit. Comparing the fitted models follows the principle of visualizing the model in the data space (Wickham et al. 2015). This app has been a valuable tool for debugging and evaluating the algorithmic behavior. By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model, particularly when projections are involved. By understanding the effect in 2-D, we aim to better understand how the algorithm might also behave in high-dimensional settings.
Figure 7: Comparing the fits for five-cluster Gaussian mixtures. Modification 2 of PPtree is used. This produces lower error and reasonable decision boundaries.
4.2 Numerical performance
@@
We compare the predictive performance for the extended PPtree algorithm with several classification methods, CART, PPtree (LDA, PDA), random forest, PPforest, and SVM, using 94 benchmark data sets. (LDA indicates PP conducted with the linear discriminant index, and PDA indicates PP conducted using the penalized linear discriminant index.) Figure 8 summarizes the performance on data sets where the modified algorithm outperformed other tree methods, and ?@tbl-data-tab2 summarizes the main characteristics of each data set, including the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, the biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables. (The Appendix contains the links to the original data sources, full performance, and data set summaries.)
@@
In the simulation, two-thirds of the observations in each data set are randomly selected for training (using stratified sampling), and the remaining one-third is used for testing and computing the prediction error. This process is repeated 200 times to assess the variance in performance.
@@
Figure 8: Average test error with 95% error bars from 200 training/test splits of benchmark data for four models: rpart, PPtree, PPtreeExt modifications 1 and 2, where the latter performs better than the others. The order of the plots is based on the pattern of errors: at top left PPtreeExt (mod 2) is best but only slightly better than rpart, and at bottom right each of the PPtree versions iteratively performs better than rpart.
Figure 8 summarises the performance for the 18 data sets, where PPtreeExt with modification 2 outperforms the other tree algorithms. In three of these data sets, it also has the lowest error than the ensemble-based classifiers, also. The plots are ordered according to the pattern of test error. In data sets data20-3, data27-a11, data27-a31, data27-b31, data27-c11, data27-c31, data37-5, data38, and data63, only PPtreeExt (mod 2) outperforms rpart. For data34 and data49-2 results are mixed. For the other data sets (data49-3, data22-1, data49-4, data56-1, data67, data69-1, data69-2), each of the methods does better than rpart. The variability across training and test splits is high for some data sets (data38, data22-1). Data sets data20-3, data49-2, and data69-1, representing a range of numerical performance, are chosen to investigate why PPtreeExt performs better, using high-dimensional visualisation in the next section.
@@
4.3 Visual explanations of performance
Exploring high dimensions is tricky, so it is helpful to set some objectives before leaping into the vast space. We would expect the modified PPtree algorithm to
  • outperform rpart when the best separation between groups is in oblique directions
  • and outperform PPtree when class variances are different, or when there are multiple clusters for any class.
This can be investigated visually using a tour. There are several strategies to use for approaching this:
@@ 4.3 Visual explanations of performance
    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart, and PPtree, as can be done using the classifly package (Wickham 2022), and then
@@
      record a sequence of projection bases that shows the boundary created by the method in focus.
@@
      use the same projection sequence to view the boundaries for the three different methods, and possibly use a slice tour to focus on the boundary near the center of the data domain.
Figure 9 shows two projections of the data20-3 illustrating why PPtreeExt performs better. The projections are 2-D from the full 5-D space. Color indicates the two classes (0 is blue and 1 is red). The circles represent the projection coefficients. Using a clock dial analogy, in the first projection (a), V3 is pointing towards 1 o’clock, and V4 is pointing towards 10. V2 is also pointing to 1, but is very small. The difference in the two groups is in the direction of 1, so V3 is likely mostly contributing to this, with a small contribution from V2. In the second projection (b) V3 points to 5, V4 points to 2, V5 points to 3, and V2 points to 9. The separation is in the direction of 5 o’clock, suggesting again that V3 is primarily important. Group 0 is split on either side of group 1, which means that the PPtree algorithm would not be able to capture both clusters, but modification 2 would. The separation between the two classes is mostly on V3 but small contributions from other variables help to improve the separation. This means that rpart will not work quite as well because it will focus on splitting V3, ignoring the other variables.
@@
Figure 9: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective. Class 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal. Secondly, the separation between the two groups is primarily V3 with small contributions from other variables, which gives PPtreeExt a slight edge over rpart.
From Figure 8, for data49-2, we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again. This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging than data20-3. A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure. In Figure 10 plot (a) shows PC 1 vs PC2, with points coloured by the two classes. PC 1 explains 99.8% of the variability in the predictors, but it explains nothing of the class structure. This is not uncommon for PCA, where an interesting structure can be found in the smaller components. Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences. Each class appears to have three linear pieces pointing in different directions. This supports why PPtreeExt might work well: in the combination of variables, the classes are nearly separable with multiple linear cuts. It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do PP with the guided tour. The optimal projection is shown in plot (c). It is computed on 26 PCs, and many of these contribute to this projection as seen by the mess of lines inside the circle representing the axes. The two classes are mostly linearly separable in this projection, but there is no clear gap between the clusters. The linear combination with so many variables is what causes some difficulties for a simple tree algorithm. The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, mean that the PPtreeExt can find better separations than PPtree.
@@
Figure 10: Examining why data49-2 is best classified using PPtreeExt. Because there are 100 predictors, for visualisation it is useful to reduce the dimension using principal component analysis, and then examine the separation between groups. PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a). PC2 and PC3 (b) have an interesting difference between the two classes, which shows why multiple linear splits might be useful. A guided projection pursuit tour on the first 26 principal components (c) shows that these classes could be separated with a linear combination. This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
From Figure 8 we see that all of the PPtree versions perform better than an rpart fit on data69-1. This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl). This is a sufficiently small-scale problem to examine using a tour, and we initially use a guided tour with the LDA index. This produced the projection shown in plot (a), where we see the primary difference between the three classes. These are each similar size, but oriented in such a way that in this 2-D projection, they form a triangle. Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively. This uses approach 2 above: created by predicting the class of points randomly generated from a uniform distribution in a 21-D cube of the same range as the data. The slice cuts through to the center of the box. We can see that PPtreeExt essentially fits the data shape neatly, but the rpart boundary is very messy. This data needs oblique cuts, which PPtreeExt provides but rpart does now. The modified algorithm handles the variance differences better than the original PPtree.
@@
Figure 11: Examining why data69-1 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt (mod 2) fitted model, and (c) boundary produced by rpart fitted model. The same projection is used in all plots to allow comparison. The boundary is examined by predicting the class of a uniform random sample of points in the data space, and viewed using a slice tour. The boundary for PPtreeExt (mod 2) mirrors what we learn from the data, that it is carving the data space into three equiangular chunks, but rpart is not capturing these linear combinations.
@@ 5 Discussion
This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt. These changes result in a more flexible classification method that leverages linear combinations of variables. The performance study showed that the predictive performance of PPtreeExt with modification 2 is better than CART and PPtree in 18 out of 94 data sets. A web app was developed and shared to help understand and debug modifications to the algorithm.
@@
Interpretability is increasingly important because predictions from algorithms may need to be defended in practice. We have shown how to examine the data and the fitted model using high-dimensional visual methods to understand why the PPtree extension performs better than other tree methods on some data sets. This approach is generally useful for assessing and diagnosing model fits.
@@
Ultimately, this may provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva et al. 2021) as available in the PPforest R package (da Silva et al. 2025). The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva et al. 2025). Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree. Using the PDA index (penalised discriminant analysis) also helps when there are many variables. Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries. This can be accomplished using the current PPforest package.
6 Supplementary Materials
This article was written in Quarto (Allaire et al. 2024), using the R package ggplot2 (Wickham 2016) for visualizations. The files to reproduce the article and results are available at https://github.com/natydasilva/PPtreeExt_paper_JCGS. Animated gifs corresponding to Figure 9, Figure 10 and Figure 11 are in the gifs folder. The code directory contains the scripts for the simulations and to explore the model fits in high dimensions.
The PPtreeExt package provides the modified PPtree algorithm. The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and the app code is provided in the PPtreeExt package, so that it can be used locally for new algorithm developments.